# When Do Transformers Learn Heuristics for Graph Connectivity?

### v1

@@ D.2 Training Dynamics of Disentangled and Standard Transformers
[…]
Figure 8: We plot the model behavior of a 3-Layer Disentangled Transformer model trained on $\mathsf{ER}(n=64)$ graphs. They also quickly pickup almost layer-wise equivariant properties (measured by Eqn. 6). All layers show very small projection error onto the $A\otimes I_{n}+B\otimes J_{n}$ decomposition, resonating our theoretical claims in Theorem 4.6.
@@
[…]
Figure 10: (left) Standard Transformers models studied in §3.3 also hit its capacity wall at $3^{L}$, showing that our theoretical results transfer beyond the theoretical simplification of disentangled transformers. (right) Standard Transformer models also learn an almost layer-wise equivariant solution measured by Eqn. 6.
@@ D.3 Scaling Effects of Diameter and Capacity
[…]
[width=]figures/restrict_diam/dt_diam_analysis/diam_2.png
@@
[…]
[width=]figures/restrict_diam/dt_diam_analysis/diam_3.png
@@
[…]
[width=]figures/restrict_diam/dt_diam_analysis/diam_4.png

### v2

@@ D.2 Additional Experiments on Disentangled and Standard Transformers
Figure 7: We plot the model behavior of a 3-Layer Disentangled Transformer model trained on $\mathsf{ER}(n=64)$ graphs. They also quickly pick up almost layer-wise equivariant properties (measured by Eqn. 3). All layers show very small projection error onto the $A\otimes I_{n}+B\otimes J_{n}$ decomposition, resonating our theoretical claims in Theorem 4.7.
@@
Figure 9: (left) Standard Transformers models studied in §3.3 also hit its capacity wall at $3^{L}$, showing that our theoretical results transfer beyond the theoretical simplification of Disentangled Transformers. (right) Standard Transformer models also learn an almost layer-wise equivariant solution measured by Eqn. 3.
In Figure 10, we show that when evaluated in $\mathsf{2Clique}$ dataset, the one trained with the right data generalize better.
Figure 10: Under the same setup as Fig. 6, when tested on $\mathsf{2Clique}$ graphs, the one trained with the right data is able to generalize better.
@@ D.3 Scaling Effects of Diameter and Capacity
Figure 11: With 1-layer Disentangled Transformers with capacity $\mathsf{Cap}=3$ following Theorem 4.5, we vary $d$ such that we restrict our training graphs to have $\operatorname{diam}(G)\leq d$. We also vary the edge probability of our training distribution $\mathsf{ER}(n=8,p=\cdot)$ for generality. We test on $\mathsf{2Chain}(n=8,k=\cdot)$ graphs with $k=2$ or 33 and show the exact match accuracy on configurations where the accuracy is non-zero for readability. We find if the training $d\leq\mathsf{Cap}$, models still learns the algorithmic solution up to problem size $d$ (see $d=2,k=2$ case on the left in orange) but fails to length generalize (see $d=2,k=3$ in orange on the right). On the other hand, if the training $d>\mathsf{Cap}$, model struggles to learn the algorithmic solution (see $d=4$ cases in red on both $k=2$ or 33). The best case overall is when setting $d=\mathsf{Cap}$, i.e., preventing the model from seeing beyond-capacity samples but still preserving at-capacity samples for better generalization. As shown in the green lines, with $d=3$, model achieves balanced testing accuracy on both $k=2$ and $3$.
@@
Appendix E Additional Related Work
Mechanistic Interpretability of Transformers.
A growing body of work reverse-engineers the algorithmic circuits that Transformers learn for tasks like copying, induction, and reasoning (Elhage et al., 2021; Olsson et al., 2022; Wang et al., 2023; Brinkmann et al., 2024). These can range from Fourier-style circuits for modular addition (Nanda et al., 2023; Zhou et al., 2024c) to Newton-like updates for in-context linear regression (Fu et al., 2024a). Researchers validate hypotheses by compiling programs into model weights (Lindner et al., 2023), decompiling models into code (Friedman et al., 2023), and using causal interventions to localize function (Chan et al., 2022; Meng et al., 2022; Yao et al., 2024; Chang et al., 2024). Theoretical work on inductive biases, like a preference for low-sensitivity functions, helps explain why models often favor robust heuristics over exact algorithms (Vasudeva et al., 2025).